# D4C glove-train: solving the RPM and Bongard-logo problem by distributing and Circumscribing concepts

### 17 Mar 2024 v3

@@ V-B D2C phase 2: The concept delineation method based on a new loss function term
Specifically, in this paper, we exclusively retain the primary representational components of all questions within the set $\{Z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which specifically correspond to $\{Z_{ij}|\,i\in[1,b],j\in[1,7]\}$. Subsequently, we randomly partition the primary representation of each case within $\{Z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{Z_{ih}|\,i\in[1,b]\}$ and $\{Z_{i\tilde{h}}|\,i\in[1,b]\}$, where $\{Z_{ih}|\,i\in[1,b]\}\cup\{Z_{i\tilde{h}}|\,i\in[1,b]\}=\{Z_{ij}|\,i\in[1,b],j\in[1,7]\}$ and $\{Z_{ih}\}\cap\{{Z}_{i\tilde{h}}\}=\emptyset$. By aggregating the vectors within $\{Z_{ih}|\,i\in[1,b]\}$ and $\{Z_{i\tilde{h}}|\,i\in[1,b]\}$, we derive the corresponding vectors $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$. Ultimately, we compute InfoNCE loss function for constraining the similarity among $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$ as the additional loss function item. The specific calculation process can be explained by:
@@
Furthermore, the spirit of D2C method can be integrated with Lico-Net by introducing images from other case statements as negative examples within the cases. And this process is illustrated in figure 8.
@@
In summary, D2C is a unpretentious and effective approach. Since the determination of image attributes in Bongard-Logo relies heavily on the concepts delineated by the main and auxiliary groups within each case, D2C tackles the problem by decoupling the concepts between different cases. However, regarding the RPM problem, which remains consistency in the identification of image attributes, this paper holds a pessimistic view on the effectiveness of D2C in solving RPM.
VI Distribution Distance-based Discriminator of Case (D3C)
VI-A D3C for Bongard-Logo
The ideal approach to address the Bongard-Logo problem involves training an encoder capable of encoding images from a Bongard-Logo case into representations, denoted as $z$. In this context, the representations of positive instances (samples in primary group) belong to the distribution $p(z)$, while the representations of negative instances (samples in primary group) belong to $q(z)$. Importantly, there should be minimal overlap between $p(z)$ and $q(z)$ in terms of cross-measure.
Specifically, this paper hypothesizes that the representation of the primary group samples in a Bongard-Logo case follows a Gaussian distribution, while the representation of the auxiliary group samples follows a mixed Gaussian distributions. We utilize an arbitrary backbone network $g(z|\,x)$ to approximate the distribution of image representations within Bongard-Logo cases by converting the process of $g(z|\,x)$ into $g(\mu,\sigma|\,x)$. The vector encoded by the backbone $g(z|\,x)$ is divided into two equal-length segments. The first half of the encoded vector is treated as the mean of a Gaussian distribution, while the second half represents the logarithm of variance, effectively normalizing the output vectors into Gaussian distributions and distributing the image representations of Bongard-Logo case. Our approach mimics the form of Variational Autoencoder (VAE). We posit that describing the image of case using distributions, rather than vectors, offers greater reasonableness.
With the estimated distributions of representations for each image in the Bongard-Logo cases, we compute the Sinkhorn distance between these distributions using reparameterization technique. Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $g(\mu,\sigma|\,x)$, aiming to train an effective discriminator. Specifically, we optimize the network by utilizing a binary cross-entropy loss function to enforce constraints on the exponential function of the negative Sinkhorn distance with base $e$ ($e^{D}\in[0,1]$, where $D\in[0,+\infty]$ is the values of Sinkhorn distance), aligning with the case’s underlying logic mentioned in DC. This process can be expressed as figure 9. The D3C Method, applied to a batch of Bongard-Logo case distributions encoded by the backbone network $g(\mu,\sigma|\,x)$, can be expressed as follows:
${\ell_{\text{D3C}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
$b$ is the batchsize of training. $S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
Figure 9: Feedforward process of D3C
VI-B D3C on Lico-Net for RPM
Notably, in contrast to the Bongard-Logo database, both RAVEN and PGM possess well-defined and comprehensive Meta data in the form of (auxiliary) rule labels. These rule descriptions facilitate the construction of a rational concept space, with non-overlapping concepts within the rule labels. This paper introduces the rule labels from the RPM database into the Lico-Net, aiming to enhance the reasoning accuracy and human-perspective rule interpretability of Lico-Net.
Specifically, under the rule description frameworks of RAVEN and PGM, there exists a many-to-one relationship between image progression patterns and rule. For instance, both single-step and double-step increments, as well as single-step and double-step decays in image attributes, are described as rule: “progressive”. Therefore, this paper decides to distribute the concept of rule from the RPM database and vectorize the image progression patterns. We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
In practice, we utilized a standard Transformer-Encoder to convert the textual descriptions of Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors termed as progression pattern vectors. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
A specific RPM case is encoded into $N\times M$ progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ from $L$ viewpoints by Lico-Net as aforementioned. To align Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net is set according to the instance of the RPM problem, depending on the number of mutually decoupled concepts present in the Meta data of that instance. For example, in PGM, there are two such concepts: shape pattern and line pattern, and $N$ is set to 2. The special setting of $N$ in Lico-Net can be viewed as $N$ slots, which can be aligned with the $N$ decoupled concepts described in the Meta data, respectively. By synthesizing all perspectives—specifically, by calculating the mean of the progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ across all $L$ viewpoints—Lico-Net encodes $N\times M$ vectors, denoted as $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$, representing the full-viewpoint progression patterns. This process can be expressed as:
$\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$
$=\frac{1}{L}\sum_{l=1}^{L}\{P_{nm}|\,n\in[1,N],m\in[1,M]\}_{l}$ (8)
The $l$ represents the index of viewpoints. The $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$ can be treated as $N$ sets of $M$ sampling results. This enables the optimization of distribution distances based on the Sinkhorn distance, thereby enhancing the overall performance. Apparently, this paper utilizes the Sinkhorn distance to constrain the distance between $M$ progression pattern vectors and their corresponding distributions’ sampling results. The $N$ decoupling concept presented in Meta data is employed to concurrently impose constraints on the $N$ sets of $M$ progression pattern vectors, and the calculate porcess can be exprecessed as:
The $\overline{P}_{nm}$ represents the full-viewpoint progressive pattern vectors encoded by Lico-Net, and $\{r_{ks}\}$ represents the $S$ sampling results from the distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of the rule description. It is evident that the specific value of $\tilde{k}$ is contingent upon the description of Meta data corresponding to the particular case of the RPM problem. symbol “$\tilde{k}|\,n$” represents that the specific value of $\tilde{k}$ is depended on $n$. This dependency realizes the fact that each slots in Lico-Net are aligned with their corresponding concepts. The $\text{Sinkhorn}(\cdot,\cdot)$ represents the Sinkhorn distance operator. The other logic of symbol in function (VI-B) has mentioned. It becomes evident that we are employing the Binary Cross Entropy (BCE) loss function to impose constraints on the exponential negative Sinkhorn distance result. A more detailed feedforward process is illustrated in the figure 10.
Figure 10: Feedforward process of Lico-Net aligned with D3C
VI-C A low-cost loss function for the D3C, the Hermes for the Zeus
In our study, we opted for cosine similarity, considering normalized vectors $u$ and $v$, where $||v||$ and $||u||$ are equal to 1, and Lipschitz constant $L=||v||=1$. It is reasonable to choose the mean of cosine similarity between sampling vectors as the distance metric function instead of Sinkhorn distance.
Correspondingly, due to the change in distance metric function, the range of function values also undergoes transformation. To be specific, the cosine similarity ranges from -1 to 1, whereas the Sinkhorn distance spans from 0 to positive infinity. Consequently, the range of values for the exponential function of the negative Sinkhorn distance with base e becomes 0 to 1. This alteration in value ranges led us to replace BCE with infoNCE as the loss function. The constraint logic of infoNCE is identical to the previously mentioned DC method. The calculate process of new method applied on Bongard-Logo can be expressed as:
${\ell_{\text{D3C-cos}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
The $t$ is set to $10^{-6}$ and $S$ is set to $5$ in function (VI-C) and (VI-C). The logic of symbols in function (VI-C) and (VI-C) has mentioned. It is evident that, compared to function (VI-A) and (VI-B), the demand for sampling frequency $S$ in D3C-cos has been significantly reduced.
The success of such simplification can be attributed, in part, to our modeling of the distribution of pattern descriptions as a Gaussian distribution. It is worth noting that the D2C and D3C methods are not in conflict. The D2C method attempts to redefine concept boundaries for reasoning problems. According to D3C, depicting concepts in the form of distributions is more reasonable compared to using vectors.
@@ VII-B High-quality negative examples lead to more rigorous concept boundaries
Our proposed D2C method has outlined a new problem boundary for deep networks through the resampling of negative examples and the introduction of a new loss function. The experimental results merely suggest that this boundary is more suitable for the scenario of missing conceptual supervision signals compared to the boundary described by the original negative examples in Bongard-Logo. However, it remains uncertain whether the new boundary is too tight or not tight enough. In other words, D2C introduces a new variance in the solution distribution for the network, and whether this variance is larger or smaller compared to the variance of the correct solution distribution is uncertain. In the context of the Bongard-Logo problem, when there is no concept supervision signal, the variance provided by D2C is closer to the variance of the correct solution distribution compared to the variance described by the original negative examples.
Clearly, D2C has specific requirements for the dataset. As expected, empirical findings indicate that this method is insufficient for defining more precise boundaries in the context of the RPM problem. Contrary to Bongard-Logo problems, the RPM problems claim that different cases share a core set of concepts, ensuring discreteness and stability among them. This shared nature renders methods like D2C unsuitable for addressing RPM problems. In other words, image attributes within RPM cases are strictly constrained by definitions of “rules” or “progressive patterns”, maintaining consistency in ascribed values across all RPM cases. The varying attributions of image properties in the Bongard-Logo problem resulting from different clustering basis among cases, which adds to problem complexity and difficulty, the inclusion of positive examples from other cases as negative examples can delineate concept boundaries between two distinct cases to some extent, adhering to a cautious approach to minimize concept infringement. However, this strategy is clearly inapplicable when dealing with RAVEN and PGM problems due to significant differences in multiple aspects.
This paper introduces a novel approach that aims to provide a tighter boundary for the discriminator by continuously generating confusing representations and challenging the discriminator. Through iterative rounds of defending against the attacks from generator, the discriminator enhances its discrimination and classification abilities, resulting in improved performance and accuracy.
VII-C discriminator for Bongard-logo
@@
Effective Bongard-Logo discriminators are abundant, with our preceding works such as PMoC and Triple-CFN demonstrating notable success. These discriminators are amenable to integration with the D2C and D3C methods proposed in this paper, and we have conducted pertinent experiments to explore their combination. These discriminators have exhibited their proficiency in accurately identifying and classifying Bongard-Logo cases, contributing significantly to the advancement of research in this domain.
@@
With regards to discriminators capable of handling adversarial tasks, the Wasserstein GAN, an excellent adversarial generative network, proposes a noteworthy requirement. The discriminator must be able to undertake a task that involves measuring the distance between samples generated by the generator and real samples. This necessitates that the discriminator must satisfy Lipschitz continuity to a certain extent.
@@
In this paper, the choice of PMoC constrainted by spectral normalization as the discriminator for the D4C method on Bongard-logo is motivated by the explicit measurement of the distance between auxiliary group sample and primiary group sample distributions provided by the PMoC approach. In contrast, the task of Triple-CFN is relatively more complex, as it not only measures the distance between primary and auxiliary samples but also considers the distance between primary samples. Therefore, compared to PMoC, this article has more design for the generator of Triple-CFN.
VII-D generator for Bongard-logo
@@
This paper argues that, instead of forcing the generator to produce samples that can confuse the discriminator, it is more effective to generate representations of samples. In terms of the intuitive metric of computational complexity, this approach reduces the pressure on the network.
@@
For the Bongard-Logo problem, this paper models the overall distribution as a mixed Gaussian distribution. Within this framework, each individual primary group in a Bongard-Logo case is modeled as a separate Gaussian distribution. Consequently, the generator of D4C designed for the Bongard-Logo problem possesses a relatively straightforward structure. The generator’s objective is to encode a distribution based on the representations of the primary group in a given Bongard-Logo case, ensuring that any representation sampled from this distribution is capable of effectively confusing the discriminator.
@@
Specifically, we use a standard Transformer-Encoder to encode a Gaussian distribution based on the primary samples’ representations in a Bongard-Logo cases, along with one optimizable vectors. The encoding approach is similar to D3C, where we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and variance logarithm of a Gaussian distribution. We then sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must exert maximum effort to counter them. The training objective of the generator is to generate representations for Bongard-Logo cases that achieve higher scores in the discriminator compared to the representations of the original negative examples in the case. However, the training objective of the discriminator is clear-cut: to classify these generated representations as negative examples. The alignment of the generator and PMoC is illustrate in figure 11.
@@
The “$f_{\theta}(z_{ij}|x_{ij})$” and “$g_{\omega}(p^{\prime}_{i}(z_{ij}|y)|\{z_{ij}\}_{j=1}^{6},z_{ij^{\prime}})$” in figure 11 is the perceptron and logic calculation module in PMoC.
For the generator of Triple-CFN, it is necessary to decay the generated representations. Specifically, when the generator with aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation needs to be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of 0.9:0.1 to decay the confusing representation vector. This is due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation, making it easy for a generator based on the Transformer-Encoder to generate realistic confusing representations, causing Triple-CFN to collapse. Therefore, this article designs the above decay process. The alignment of the generator and Triple-CFN is illustrate in figure 12. The “$f_{\theta}(z_{ij}|x_{ij})$” in figure 12 is the encoder in Triple-CFN.
@@
VII-E discriminator of RPM
In terms of discriminating RPM problems, this paper places significant faith in the Lico-Net model as an exceptional baseline. The spectral normalization is anticipated to elevate Lico-Net’s performance, making it a highly suitable discriminator for the D4C method.
VII-F generator of RPM
Distinct from the Bongard-Logo problem, the answer for a single case in the RPM problem can be represented as a mixed Gaussian distributions due to its non-uniqueness. Thus, Our aspiration is for the generator, tailored for the D4C method within the RPM context, to encode a distribution that represents a mixed Gaussian distribution. This paper employs a glove-shaped network to simulate the process of sampling from a mixed Gaussian distributions. Specifically, we utilize a standard Transformer-Encoder to process the representations of all the problem images in an RPM case, along with $W$ optimizable vectors (where $W$ is a hyperparameter.) and a stochastic promoter (these $W+1$ vectors, when arranged together, surprisingly resemble the shape of a glove). The promoter is a vector sampled from a multivariate Gaussian distribution, where the mean of the distribution is optimizable while the variance is fixed at 1.
It is worth noting that we have designed a novel positional encoding embedding method for the backbone of our RPM generator, the Transformer-Encoder, to restore the $3\times 3$ image arrangement of the RPM problem matrix. For the PGM problem, our positional encoding scheme involves assigning unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case. The positional encodings for images in the lower triangular region are correspondingly symmetrical with respect to the diagonal, mirroring the encodings of their counterparts in the upper triangular region. However, this paper does not demonstrate the superiority of D4C on the relatively less challenging RAVEN dataset.
Back to the main topic. After obtaining the attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from the attention results and process them into $W$ distributions using a linear layer like aforementioned. Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard-gumbel-softmax, yielding highlighted probabilities for the $W$ distributions. We then sample representations as confusing representations from the highlighted distributions each time. In the technical implementation, we employed the reparameterization trick to draw samples from each distribution, yielding $W$ sampling results. Subsequently, we performed a weighted sum of these sampling results using the hard probabilities outputted by the hard-gumbel-softmax, thus completing the entire mixed-Gaussian sampling process. The approach for modeling and training the network, as presented in this paper, is coined as “D4C glove-train”. A more elaborate illustration of the forward process is provided in the accompanying figure 13.
Figure 13: Feedforward process of D4C on RPM problem

### 25 Mar 2024 v4

@@ VI-A2 D2C phase 2: The concept delineation method based on a new loss function term
Specifically, in this paper, we exclusively retain the representation of primary group within the set $\{z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which specifically correspond to $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$. Subsequently, we randomly partition the primary representation of each case within $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, where $\{z_{ih}|\,i\in[1,b]\}\cup\{z_{i\tilde{h}}|\,i\in[1,b]\}=\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ and $\{z_{ih}\}\cap\{{z}_{i\tilde{h}}\}=\emptyset$. By taking the mean of the vectors within $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, we derive the corresponding vectors $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$. Ultimately, we compute InfoNCE loss function for constraining the similarity among $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$ as the additional loss function item. The specific calculation process can be explained by:
@@
In essence, the phase 2 of D2C attempts to achieve a similar goal as its first phase by formulating a new loss function, approaching the task from a different angle. Nevertheless, this strategy renders the phase 2 of D2C primarily compatible with backbone that can be abstracted as $f(z|x)$. This adaptability enables a majority of image encoders, such as ResNet, ViT, and Triple-CFN[32], to be compatible with the complete D2C methodology. However, the application of PMoC[31], which relies on probability-based measures to assess image identity, is restricted to the first phase of D2C.
VI-B D2C on RPM probelm
Furthermore, the spirit of D2C method can be integrated with Lico-Net by introducing images from other case statements as negative examples within the cases. However, the application of D2C in RPM discriminators is limited to its phase 1, as it becomes challenging to forcibly apply the phase 2 to these discriminators. This process is illustrated in figure 15.
@@ VI-B D2C on RPM probelm
In summary, D2C is a straightforward approach. In the context of Bongard-Logo, where the determination of image attributes heavily relies on the concepts outlined by the primary and auxiliary groups within each case, D2C addresses this challenge by decoupling the concepts across different cases. Nevertheless, regarding the RPM problem, which emphasizes consistency in image attribute identification, this study maintains a pessimistic outlook on the efficacy of D2C in tackling RPM. It is worth emphasizing that the D2C and D3C methodologies are not mutually exclusive. Specifically, D2C endeavors to redefine concept boundaries for reasoning problems. According to D3C, representing concepts in the form of distributions is a more plausible approach compared to utilizing vectors.
@@ VII-B High-quality negative examples lead to more rigorous concept boundaries
The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative instances and incorporating a fresh loss function. Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of conceptual supervision signals, as compared to the boundaries delineated by the original negative instances in Bongard-Logo. Nonetheless, the precise tightness of the new boundary remains indeterminate.
In essence, D2C introduces a new variance to the solution distribution of the network, the magnitude of which relative to the variance of the correct solution distribution is indeterminate. In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of conceptual supervision signals, as compared to the variance described by the original negative instances.
Evidently, the efficacy of D2C is contingent upon specific dataset attributes. As anticipated, empirical evidence suggests that this approach may be insufficient for delineating precise boundaries within the framework of the RPM problem. Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct instances, ensuring discreteness and stability. This shared nature precludes the utility of methods like D2C in addressing RPM challenges. Specifically, image properties within RPM instances are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all instances.
The Bongard-Logo problem exhibits varying attributions of image properties due to differing clustering bases among instances, thereby augmenting problem complexity and difficulty. Incorporating positive instances from other cases as negative examples can, to some extent, delineate conceptual boundaries between distinct cases, adhering to a cautious approach aimed at minimizing conceptual infringement. However, this strategy proves inapplicable in addressing challenges presented by RAVEN and PGM due to substantial disparities across multiple dimensions.
In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative instances within Bongard-Logo cases. Conversely, its limitations in tackling the RPM problem can also be ascribed to its insignificant improvement in the quality of negative examples within RPM scenarios. To address the limitations inherent in the D2C approach, this paper introduces the D4C (Defense-Driven Discrimination of Distributed Concepts) method. The underlying idea of D4C is to generate higher-quality negative instances and samples within abstract reasoning problems. This approach aims to delineate a more appropriate problem boundary for the discriminator as compared to the original problem setup in abstract reasoning. By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of conceptual boundaries and foster further advancements in the realm of abstract reasoning.
VII-C Impelmention of D4C method
The D4C methodology proposed in this paper emphasizes on devising a more suitable concept boundary for both RPM and Bongard-Logo discriminators, aiming to enhance their problem-solving capabilities. To achieve this objective, a distinctive approach centered on the generation of confusing samples is introduced. These samples are intentionally designed to challenge and improve the discriminator’s accuracy. A dedicated generator has been devised to carry out this generation process. Through iterative defense against attacks launched by this generator, the discriminator progressively refines its discrimination and classification skills, leading to overall performance enhancement in both areas. Specifically, within this study, a generator has been tailored to augment the capabilities of two distinct Bongard-Logo discriminators: PMoC and Triple-CFN. Additionally, this generator can also be directly utilized for the refinement of Lico-Net. Nevertheless, acknowledging the inherent disparities between RPM problems and Bongard-Logo challenges, additional designs have been incorporated into the generator to cater more effectively to the refinement requirements of Lico-Net.
In this section, we delve into the design of the discriminator and generator within the D4C framework as it pertains to the Bongard-Logo and RPM problems.
VII-C1 discriminator for Bongard-logo
@@ VII-C Impelmention of D4C method
Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC and Triple-CFN, achieving notable successes. These discriminators seamlessly integrate with the D2C and D3C methodologies proposed in this study, and pertinent experiments have been conducted to explore their combinations. These discriminators excel in accurately identifying and classifying Bongard-Logo instances, significantly contributing to the advancement of research in this domain.
@@
Regarding discriminators capable of handling adversarial tasks, the Wasserstein GAN, a prominent adversarial generative network, poses a noteworthy requirement. The discriminator must undertake the task of measuring the distance between samples generated by the generator and real samples. This necessitates that the discriminator satisfies Lipschitz continuity to a certain extent.
@@
In this study, the choice of PMoC, constrained by spectral normalization, as the discriminator for the D4C method on Bongard-logo is motivated by its explicit measurement of the distance between auxiliary and primary group sample distributions. Conversely, the task of Triple-CFN is relatively more intricate as it considers not only the distance between primary and auxiliary samples but also the distance among primary samples. Consequently, compared to PMoC, this study entails more elaborate designs for the generator of Triple-CFN.
VII-C2 generator for Bongard-logo
@@
In this paper, we argue that instead of forcing the generator to produce samples that can confuse the discriminator, it is more effective to generate representations of confusing samples, referred to as confusing representations. This approach reduces the computational burden on the network, as measured by intuitive metrics such as computational complexity.
@@
For the Bongard-Logo problem, we model the overall distribution as a mixture of Gaussian distributions. Within this framework, each individual primary group in a Bongard-Logo case is represented as a separate Gaussian distribution. Consequently, the D4C generator designed for the Bongard-Logo problem exhibits a relatively straightforward structure. The generator aims to encode a distribution based on the representations of the primary groups in a given Bongard-Logo case, ensuring that any representation sampled from this distribution is capable of effectively confusing the discriminator.
@@
Specifically, we employ a standard Transformer encoder to encode a Gaussian distribution based on the representations of the primary samples in a Bongard-Logo case, along with an optimizable vector. This encoding approach bears similarity to D3C, wherein we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and logarithm of variance for the Gaussian distribution. Subsequently, we sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must exert maximum effort to counter them. The generator’s training objective is to generate representations for Bongard-Logo cases that achieve higher scores in the discriminator compared to the representations of the original negative examples in the case. Conversely, the discriminator’s training objective is unambiguous: to classify these generated representations as negative examples. The aforementioned approach can be seamlessly applied to the PMoC model, providing a direct and effective means of enhancing its performance. The alignment between the generator and PMoC is illustrated in Figure 16.
@@
In this figure, the “$f_{\theta}(z_{ij}|x_{ij})$” and “$g_{\omega}(p^{\prime}_{i}(z_{ij}|y)|\{z_{ij}\}_{j=1}^{6},z_{ij^{\prime}})$” represent the perceptron and logic calculation modules in PMoC, respectively[31].
For the generator in Triple-CFN, it is essential to attenuate the generated representations. Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to attenuate the confusing representation vector. This attenuation process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation. This makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse. Therefore, we have designed the aforementioned attenuation process. The alignment between the generator and Triple-CFN is illustrated in Figure 17. In this figure, “$f_{\theta}(z_{ij}|x_{ij})$” represents the encoder in Triple-CFN.
@@
VII-C3 discriminator of RPM
In addressing Raven’s Progressive Matrices (RPM) problems, this study places significant confidence in the Lico-Net model, considering it as an exceptional benchmark. It is anticipated that the application of spectral normalization will enhance the performance of Lico-Net, positioning it as a highly suitable discriminator for the D4C methodology.
However, Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix. To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations. Any inconsistencies or irregularities identified within these representations will be flagged as negative instances, indicating incorrect responses. Through this rigorous process, Lico-Net strives to accurately identify and reject perplexing representations, thereby augmenting its overall precision and reliability.
VII-C4 generator of RPM
Unlike the Bongard-Logo problem, the answer to a single case in the RPM problem exhibits non-uniqueness, and thus can be represented as a mixture of Gaussian distributions. Therefore, our aspiration is to tailor a generator for the D4C method within the RPM context, capable of encoding a distribution that reflects this mixed Gaussian nature. In this paper, we employ a glove-shaped network to simulate the process of sampling from such mixed Gaussian distributions.
Specifically, we utilize a standard Transformer-Encoder to process the representations of all problem images in an RPM case, alongside $W$ optimizable vectors (where $W$ is a hyperparameter) and a stochastic promoter. Notably, when these $W+1$ vectors are arranged together, they resemble the shape of a glove. The promoter is a vector sampled from a multivariate Gaussian distribution, wherein the mean is optimizable while the variance is fixed at 1.
It is worth noting that we have devised a novel positional encoding embedding approach for the Transformer-Encoder, which serves as the backbone of our RPM generator. This encoding aims to restore the $3\times 3$ image arrangement characteristic of the RPM problem matrix. For the PGM (Progressive Graph Matching) problem, our positional encoding scheme assigns unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case. Correspondingly, the positional encodings for images in the lower triangular region are symmetrical with respect to the diagonal, mirroring the encodings of their counterparts in the upper triangular region. However, it should be acknowledged that this paper does not demonstrate the superiority of D4C on the relatively less challenging RAVEN dataset.
Moving on to the main topic, after obtaining attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from these results and process them into $W$ distributions using a linear layer, as mentioned earlier. Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard-gumbel-softmax activation, yielding highlighted probabilities for the $W$ distributions. Subsequently, we sample representations, referred to as confusing representations, from the highlighted distributions each time. In the technical implementation, we employ the reparameterization trick to draw samples from each distribution, resulting in $W$ sampling outcomes. We then perform a weighted sum of these sampling results using the hard-highlighted probabilities processed by the hard-gumbel-softmax, thus completing the entire mixed-Gaussian sampling process. The methodology outlined in this paper for modeling and training the network is coined as ”D4C glove-train”. A more detailed illustration of the forward process is provided in the accompanying figure referenced as 18.
The training objective of the glove-shaped generator remains consistent: to produce confusing representations that achieve higher scores in the discriminator compared to the original negative examples provided by the dataset. Conversely, the discriminator aims to resist the interference posed by these confusing representations.
Figure 18: Feedforward process of D4C on RPM problem